Sales Mix and Break-Even Sales

Dragon Sports Inc. manufactures and sells two products, baseball bats and baseball gloves. The Costs that tend to remain the same in amount, regardless of variations in the level of activity.fixed costs are $620,000, and the sales mix is 40% bats and 60% gloves. The unit selling price and the unit Costs that vary in total dollar amount as the level of activity changes.variable cost for each product are as follows:

Products Unit Selling Price Unit Variable Cost
Bats $90 $50
Gloves 105 65

a. Compute the break-even sales (units) for the overall enterprise product, E.
units

b. How many units of each product, baseball bats and baseball gloves, would be sold at the The level of business operations at which revenues and expired costs are equal.break-even point?

Baseball bats units
Baseball gloves units
Answer #1

a) Weighted average contribution margin = (40*40%+40*60%) = 40

Break even unit = 620000/40 = 15500 Units

b) Baseball bats = 15500*40% = 6200 Units

Baseball gloves = 15500*60% = 9300 units